OpenStudy (anonymous):

A ball is thrown upward. What is its initial vertical speed? The acceleration of gravity is 9.8 m/s^2 and maximum height is 4.3 m. Neglect air resistance. Answer in units of m/s.

OpenStudy (anonymous):

I thought I could use final velocity = initial velocity + acceleration*time, and make time=0 because it's the initial, but I don't have final velocity, I have distance instead. Is distance this case the same as displacement, and if so, is there an equation that incorporates my two known variables (acceleration and height)?

OpenStudy (anonymous):

Oh god First use s= 1/2gt*2 to find time and then use your equation v=u+at Or else u can directly use the equation v*2=u*2+2as There you go (:

OpenStudy (anonymous):

Just making sure I'm doing this right (bear with me this is my first physics class of my entire life), you should get 4.3=(1/2)(9.8)t^2, and t=.87755 seconds. So then 0 = u + (9.8)(.87755)^2, so u = -8.6??

OpenStudy (anonymous):

I found another way to do it. You can directly use v*2=u*2+2x9.81x4.3. V*2 will be 0 so don't take it in the equation. Shift 2x9.81x4.3. Towards the right, so now the equation will be 2x9.81x4.3= u*2 84.37= u*2 now shift the*2 making it the square root of 84.37 giving you 9.19= u(initial velocity) there you go:)

OpenStudy (anonymous):

Ah, thank you. That makes a lot of sense!